# PlasmoData.jl – A Julia Framework for Modeling and Analyzing Complex Data as Graphs

### v1

@@ Abstract
Datasets encountered in scientific and engineering applications appear in complex formats (e.g., images, multivariate time series, molecules, video, text strings, networks). Graph theory provides a unifying framework to model such datasets and enables the use of powerful tools that can help analyze, visualize, and extract value from data. In this work, we present PlasmoData.jl, an open-source, Julia framework that uses concepts of graph theory to facilitate the modeling and analysis of complex datasets. The core of our framework is a general data modeling abstraction, which we call a DataGraph. We show how the abstraction and software implementation can be used to represent diverse data objects as graphs and to enable the use of tools from topology, graph theory, and machine learning (e.g., graph neural networks) to conduct a variety of tasks. We illustrate the versatility of the framework by using real datasets: i) an image classification problem using topological data analysis to extract features from the graph model to train machine learning models; ii) a disease outbreak problem where we model multivariate time series as graphs to detect abnormal events; and iii) a technology pathway analysis problem where we highlight how we can use graphs to navigate connectivity. Our discussion also highlights how PlasmoData.jl leverages native Julia capabilities to enable compact syntax, scalable computations, and interfaces with diverse packages.
@@ 1 Introduction
Data appears in complex formats that require the use of advanced tools for its representation and processing. In this context, it is important to highlight that any data object (e.g., image, video, audio signal, a text string, time series) needs to be represented as a mathematical model (a mathematical abstraction) to enable processing and analysis. Examples of mathematical models that are available to do this include those from linear algebra (e.g., data is modeled as vectors, matrices, tensors), graph/network theory (e.g., data is modeled as graphs or hypergraphs), statistics (e.g., data is modeled as random variables), and topology/geometry (e.g., data is modeled as graphs, manifolds, and simplicial complexes). Understanding how to model data and to recognize alternative representations of data is key in facilitating its analysis and in extracting information/value [1, 2, 3]. For example, an image can be represented as a matrix (each entry is a pixel), as a graph (each node is a pixel and edges represent connectivity of neighboring pixels), or as a function/manifold (the additional dimension represents the light intensity). These various representations reveal different (and often complementary) aspects/features of the data object; for instance, a matrix representation can help reveal correlation structures, while a graph representation can reveal connected structures. Importantly, how to represent a dataset is a modeling decision that can impact the information extracted from the data (can influence the value of the dataset); as such, it is necessary to have proper tools that can help experiment with different data models and associated analysis tools.
@@
A modeling abstraction that has gained increased popularity in data science are graphs (networks). Graphs provide a unifying framework to capture diverse data objects because they have intuitive interpretation and they enable the use of powerful processing techniques. Simply stated, a graph is a mathematical model that comprises a set of nodes that are connected via edges; the nodes and edges are abstract objects that can encode data that may appear in different forms. In a typical graph, data attached to nodes and edges are scalar values or vectors, but more advanced graph abstractions can attach abstract objects such as algebraic models, text, or other graphs [4, 5]. Graph representations have been used in a wide variety of applications such as chemical processes [6], biological systems [7], brain networks [8, 9], hydrology [10, 11, 12], disease transmission [13, 14], and social networks [15]. An application that has recently gained substantial attention is the representation of molecules as graphs [16, 17, 18]; here, atoms are represented by nodes, bonds are represented by edges, and nodes/edges can embed diverse attributes such as atom type and bond length (represented as vectors). These representations have been used to predict diverse properties of molecules and to predict potential reactions that can occur between them.
@@
The goal of this work is to provide a general framework for modeling data as graphs. Recently, we have developed a framework for modeling optimization problems as graphs[4]; in this abstraction, nodes and edges can embed algebraic optimization problems. This abstraction has been shown to unify a wide range of structures found in optimization (e.g., optimal control, stochastic optimization, and network optimization) and this unification has enabled a number of advances in theory and algorithms. Moreover, the abstraction has facilitated access to a broad range of graph analysis and software tools that facilitate visualization and processing (e.g., graph partitioning and aggregation). Our work aims to expand this graph abstraction to model general data and with this unify data objects found across domains and leverage the use of tools from such domans. Our abstraction is implemented as an open-source Julia package that we call PlasmoData.jl. This package has been designed to readily represent diverse data objects as graphs (e.g., images, matrices, and tensors as node-weighted graphs and symmetric matrices as edge-weighted graphs) and to store data within a user-defined graph structure. Our modeling framework interfaces to diverse software packages that enables processing of graph data objects (e.g., via filtration, partition, and aggregation), facilitates the computation of descriptors using tools from graph theory and topology, and enables the use of machine learning tools (e.g., graph neural networks-GNNs). Our software design principle is analogous to those of algebraic modeling languages (e.g., JuMP or Plasmo.jl), which provide provide interfaces to packages that process/solve such models.
The goal behind the implementation of PlasmoData.jl is thus to change the way the user thinks about data (as a modeling task rather than just an analysis task). We believe that the focus on modeling can bring significant benefits in the way we explore alternative data representations that are suitable for applications, in selecting suitable tools to enable data processing, and in interpreting analysis results. We demonstrate the versatility of PlasmoData.jl and of its unifying modeling abstraction by using applications that appear in quite distinct application domains. A graphical (pun intended) abstract of this work is shown in Figure 1 and highlights how a variety of data can be modeled within PlasmoData.jl, which in turn provides access to diverse analysis tools and techniques.
@@
Figure 1: Overview of the PlasmoData.jl modeling framework. Diverse data objects can be modeled under the DataGraph abstraction; this provides access to a diverse analysis tools and techniques for analyzing the data.
@@ 2 Graph Abstraction and Software Implementation
This section introduces the DataGraph modeling abstraction along with its implementation in PlasmoData.jl, available at https://github.com/zavalab/PlasmoData.jl. We introduce a mathematical formulation for the DataGraph and discuss how it can be applied to a variety of common data structures. We also provide brief code snippets showing how these representations are implemented in PlasmoData.jl.
@@ 2.1 Mathematical Formulation
A DataGraph object is fully defined by its nodes, edges, node data, edge data, graph data, node attributes, edge attributes, graph attributes, and an adjacency matrix $A$ that encodes the graph connectivity. Importantly, the adjacency matrix also encodes the type of edges in the graph (directed or undirected).
@@ 2.2 Software Implementation
Figure 2: Visualization of the DataGraph object. The PlasmoData.DataGraph contains fundamental attributes (g, nodes, edges, node_map, edge_map, node_data, edge_data, graph_data). Each node and edge can embed data of a variety of forms, such as scalar values, vectors, or other graphs.
@@
PlasmoData.jl implements the DataGraph abstraction as shown above, and it provides a user-friendly interface and suite of functions. The code in Code Snippet 1 illustrates the basic functionality of PlasmoData.jl.
Code Snippet 1: Basic functionality for building a PlasmoData.DataGraph in PlasmoData.jl
1using PlasmoData, DataGraphPlots
@@
38DataGraphPlots.plot_graph(dg; xdim = 400, ydim = 400)
Figure 3: Visualization of the PlasmoData.DataGraph associated with Code Snippet 1.
An empty PlasmoData.DataGraph is instantiated on Line 3. The data types within the { } define the types of data that can be defined in the node, edge, and graph data (see Figure 2). In this example, we use Any for the data type in the node, edge, and graph data which will be important when adding data to this graph. Alternatively, a user can call DataGraph() without the information in { }, and the default data types in the node, edge, and graph data will be Float64. Once the PlasmoData.DataGraph object is instantiated, nodes, edges, and data can be added to the graph.
Nodes can be added to the PlasmoData.DataGraph by calling the add_node! function (Lines 6 - 10). PlasmoData.jl allows for any Julia object to be defined as a node. For example, as shown in the code snippet, the node names can be integers, strings, or symbols. This can be important, as will be shown later when representing matrices as PlasmoData.DataGraphs.
Edges can be added to the PlasmoData.DataGraph by calling the add_edge! function (Lines 14 - 18). This function takes pairs of node names as arguments which represent the edge between those two nodes. These pairs of nodes can be passed as separate arguments (Lines 14 - 16) or as a tuple of two node names (Lines 17 - 18). This progressively builds the adjacency matrix of the graph object.
Data is embedded on the graph through API functions. Data is added to nodes through the add_node_data! function (Lines 21 - 25). This function takes the following arguments: i) the PlasmoData.DataGraph, ii) the node name, iii) the node data, and iv) the attribute name for the data (must be a string). The node data is restricted to be the type defined on the PlasmoData.DataGraph (in this case, the PlasmoData.DataGraph was instantiated with type Any, so any data can be passed). As seen in the code snippet, data have different formats, including floats, strings, vectors, matrices, or another PlasmoData.DataGraph. Data is added to edges through the add_edge_data! function (Lines 29 - 33). This function behaves similarly to the add_node_data! function, but in place of the node name argument, a pair of nodes (two separate arguments) or a tuple of two node names (one argument) is supplied. Data is added to the graph by the add_graph_data! function (Line 36). This function only takes the graph data and the graph attribute, as shown in the snippet. Additional node, edge, or graph data can be defined on the DataGraph by passing the data with a different attribute name. There is no limit to the number of attributes that can be added to the nodes, edges, or graph.
One of the benefits of graph representations is that they facilitate visualization of structure. For basic plotting capabilities, we have developed an accompanying package called DataGraphPlots.jl (available at https://github.com/zavalab/DataGraphPlots.jl) that interfaces to PlasmoData.jl and that enables visualizations of PlasmoData.DataGraph objects. This package provides basic functionalities for plotting both PlasmoData.DataGraphs and PlasmoData.DataDiGraphs primarily through the Julia packages Plots.jl and GraphMakie.jl. The visualization in Figure 3 is created by calling the DataGraphPlots.jl function plot_graph (Line 38).
@@ 2.3 Modeling Common Data Objects as DataGraphs
With the general mathematical formulation for the DataGraph defined, we next outline how specific data objects (matrices, symmetric matrices, tensors) can be represented using the graph abstraction and how these can be implemented in PlasmoData.jl.
@@ 2.3.1 Matrices as a Node-Weighted Graphs
We can represent a matrix $M\in\mathbb{R}^{p\times q}$ (where $p$ and $q$ are positive integers) as a node-weighted graph with a mesh structure [19, 34], and we can express the resulting graph using the DataGraph abstraction. We use the notation $m_{i,j}\in\mathbb{R}$ for the matrix entry at row $i$ and column $j$. For convenience, we will define the integer set $\mathbb{N}_{k}:=\{1,2,...,k\}$. To represent the matrix as a graph, we define a node, $n_{*}$, for every matrix entry and place edges between all adjacent entries of the matrix. On each node, $m_{i,j}$ represents the node data. As there is a single weight in each node, we define the attribute set as $\mathcal{A}_{\mathcal{N}}=\{a\}$ (a singleton).
Any matrix can be represented with the DataGraph abstraction as:
@@
An alternative structure can also be constructed by also adding edges between diagonal elements of the matrix [35]. In this case, the set of edges becomes =$\mathcal{E}=\{(n_{i,j},n_{i+1,j}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q}\}\cup\{(n_{i,j},n_{i,j+1}):i\in\mathbb{N}_{p},j\in\mathbb{N}_{q-1}\}\cup\{(n_{i,j},n_{i+1,j+1}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q-1}\}\cup\{(n_{i,j+1},n_{i+1,j}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q-1}\}$.
PlasmoData.jl facilitates the the representation of matrices as graphs. Code Snippet 2 shows how a PlasmoData.DataGraph object can be constructed automatically by calling the matrix_to_graph function. Any matrix containing real numbers can be passed to this function. The boolean keyword argument diagonal can be passed to identify whether the diagonal edges should be included in the graph. The resulting graphs can also be visualized using DataGraphPlots.jl. In Line 10, nodes are given a fixed position (stored in the node data of the PlasmoData.DataGraph) for visualization. In Lines 13 - 20, the graph containing diagonals is plotted, where the weight data from the original matrix is used for coloring the nodes (see Line 18). Figure 4 visualizes the alternative representations. Because matrices themselves are a general abstraction for representing data, PlasmoData.DataGraph can be applied to many application. For instance, multivariate time series and grayscale images can be represented as matrices (and thus as graphs) [36, 37, 38].
Code Snippet 2: Representing a matrix as a PlasmoData.DataGraph
1using PlasmoData, Random, DataGraphPlots
@@
3random_matrix = rand(12, 12)
4
5# Convert matrix to node-weighted graph
6matrix_graph_diags = matrix_to_graph(random_matrix; diagonal = true)
7matrix_graph_no_diags = matrix_to_graph(random_matrix; diagonal = false)
8
9# Fix the node positions
10set_matrix_node_positions!(matrix_graph_diags, random_matrix)
11
12# Plot the graph
13plot_graph(
@@
20)
Figure 4: Visualization of Code Snippet 2 where a PlasmoData.DataGraph is formed from a random matrix.
2.3.2 Tensors as Node-Weighted Graphs
Any tensor $T\in\mathbb{R}^{p\times q\times r}$ (where $p$, $q$, and $r$ are positive integers) can be represented as a node-weighted graph. There are at least a couple of methods for how to represent $T$. The first approach is to construct a node-weighted mesh following the description outlined in Section 2.3.1 based on the first couple of dimensions of $T$ (i.e., the mesh structure will be $p\times q$). The third dimension is then represented as a vector of weights attached to every element of the mesh (with $r=|\mathcal{A}_{\mathcal{N}}|$). This can be used, for instance, to represent color images (which have different color channels). In addition, we use the notation $t_{i,j,k}$ for the tensor entry, where $i,j,$ and $k$ are the entries in the first, second, and third dimensions, respectively. Under the DataGraph abstraction, this has the form:
@@ 2.3.2 Tensors as Node-Weighted Graphs
PlasmoData.jl facilitates both of the representations discussed. In code snippet 3, the random tensor can be formed into the mesh structure discussed in Section 2.3.1 by calling the matrix_to_graph function (Line 6). This function recognizes that this is a 3D array and creates the mesh structure based on the first two dimensions. The third dimension are the weights on each node. The default name for these weights is the string "weight" with a number (e.g., "weight1", "weight2"), but a user can define their own names for the attributes in $\mathcal{A}_{\mathcal{N}}$. Alternatively, the user can call the function tensor_to_graph (Line 9) which performs the second method discussed above. These graphs can be visualized using DataGraphPlots.jl as before (Lines 12 - 19). In this case, the node locations are determined automatically within DataGraphPlots.jl through the NetworkLayout.jl package. A visualization of the methods in the code snippet is presented in Figure 5.
Code Snippet 3: Representing a tensor as a PlasmoData.DataGraph using a couple of different approaches
1using PlasmoData, Random, DataGraphPlots
@@
3random_tensor = rand(4, 5, 6)
4
5# Convert the tensor to a node weighted graph with 4 x 5 nodes
6tensor_graph_2d = matrix_to_graph(random_tensor; diagonal = true)
7
8# Convert the tensor to a node weighted graph 4 x 5 x 6 nodes
9tensor_graph_3d = tensor_to_graph(random_tensor)
10
11# Plot the graph
12plot_graph(
@@
19)
Figure 5: Visualization of Code Snippet 3 where a PlasmoData.DataGraph is formed from a random tensor.
@@
It is worth emphasizing that a benefit of representing matrices and tensors as graphs is that graphs are modeling objects that do not live in a Euclidean space. As such, all that matters from a graph perspective is connectivity (not location of nodes and edges in space). As such, a graph object is not affected by rotations (matrices are). This property is quite useful when processing images/video and also when processing data objects from molecular simulations (in which nodes mode randomly in space) [39].

### v2

@@ Abstract
Datasets encountered in scientific and engineering applications appear in complex formats (e.g., images, multivariate time series, molecules, video, text strings, networks). Graph theory provides a unifying framework to model such datasets and enables the use of powerful tools that can help analyze, visualize, and extract value from data. In this work, we present PlasmoData.jl, an open-source, Julia framework that uses concepts of graph theory to facilitate the modeling and analysis of complex datasets. The core of our framework is a general data modeling abstraction, which we call a DataGraph. We show how the abstraction and software implementation can be used to represent diverse data objects as graphs and to enable the use of tools from topology, graph theory, and machine learning (e.g., graph neural networks) to conduct a variety of tasks. We illustrate the versatility of the framework by using real datasets: i) an image classification problem using topological data analysis to extract features from the graph model to train machine learning models; ii) a disease outbreak problem where we model multivariate time series as graphs to detect abnormal events; and iii) a technology pathway analysis problem where we highlight how we can use graphs to navigate connectivity. Our discussion also highlights how PlasmoData.jl leverages native Julia capabilities to enable compact syntax, scalable computations, and interfaces with diverse packages. Overall, we show that the DataGraph abstraction and PlasmoData.jl Julia package are able to model data within graphs and enable useful analysis.
@@ 1 Introduction
Data appears in complex formats that require the use of advanced tools for its representation and processing. In this context, it is important to highlight that any data object (e.g., image, video, audio signal, a text string, time series) needs to be represented as a mathematical model (a mathematical abstraction) to enable processing and analysis. Examples of mathematical models that are available to do this include those from linear algebra (e.g., data is modeled as vectors, matrices, higher-order tensors), graph/network theory (e.g., data is modeled as graphs or hypergraphs), statistics (e.g., data is modeled as random variables), and topology/geometry (e.g., data is modeled as graphs, manifolds, and simplicial complexes). Understanding how to model data and to recognize alternative representations of data is key in facilitating its analysis and in extracting information/value [1, 2, 3]. For example, an image can be represented as a matrix (each entry is a pixel), as a graph (each node is a pixel and edges represent connectivity of neighboring pixels), or as a function/manifold (the additional dimension represents the light intensity). These various representations reveal different (and often complementary) aspects/features of the data object; for instance, a matrix representation can help reveal correlation structures, while a graph representation can reveal connected structures. Importantly, how to represent a dataset is a modeling decision that can impact the information extracted from the data (can influence the value of the dataset); as such, it is necessary to have proper tools that can help experiment with different data models and associated analysis tools.
@@
A modeling abstraction that has gained increased popularity in data science is graphs (networks). Graphs provide a unifying framework to capture diverse data objects because they have intuitive interpretation and they enable the use of powerful processing techniques. Simply stated, a graph is a mathematical model that comprises a set of nodes that are connected via edges; the nodes and edges are abstract objects that can encode data that may appear in different forms. In a typical graph, data attached to nodes and edges are scalar values or vectors, but more advanced graph abstractions can attach abstract objects such as algebraic models, text, or other graphs [4, 5]. Graph representations have been used in a wide variety of applications such as chemical processes [6], biological systems [7], brain networks [8, 9], hydrology [10, 11, 12], disease transmission [13, 14], and social networks [15]. An application that has recently gained substantial attention is the representation of molecules as graphs [16, 17, 18]; here, atoms are represented by nodes, bonds are represented by edges, and nodes/edges can embed diverse attributes such as atom type and bond length (represented as vectors). These representations have been used to predict diverse properties of molecules and to predict potential reactions that can occur between them.
@@
The goal of this work is to provide a general framework for modeling data as graphs. Recently, we have developed a framework for modeling optimization problems as graphs[4]; in this abstraction, nodes and edges can embed algebraic optimization problems. This abstraction has been shown to unify a wide range of structures found in optimization (e.g., optimal control, stochastic optimization, and network optimization) and this unification has enabled a number of advances in theory and algorithms. Moreover, the abstraction has facilitated access to a broad range of graph analysis and software tools that facilitate visualization and processing (e.g., graph partitioning and aggregation). Our work aims to expand this graph abstraction to model general data and with this unify data objects found across domains and leverage the use of tools from such domains. Our abstraction is implemented as an open-source Julia package that we call PlasmoData.jl. This package has been designed to readily represent diverse data objects as graphs (e.g., images, matrices, and tensors as node-weighted graphs and symmetric matrices as edge-weighted graphs) and to store data within a user-defined graph structure. Our modeling framework interfaces to diverse software packages that enables processing of graph data objects (e.g., via filtration, partition, and aggregation), facilitates the computation of descriptors using tools from graph theory and topology, and enables the use of machine learning tools (e.g., graph neural networks or GNNs). Our software design principle is analogous to those of algebraic modeling languages (e.g., JuMP or Plasmo.jl), which provide provide interfaces to packages that process/solve such models.
The goal behind the implementation of PlasmoData.jl is thus to change the way the user thinks about data (as a modeling task rather than just an analysis task). We believe that the focus on modeling can bring significant benefits in the way we explore alternative data representations that are suitable for applications, in selecting suitable tools to enable data processing, and in interpreting analysis results. We demonstrate the versatility of PlasmoData.jl and of its unifying modeling abstraction by using applications that appear in quite distinct application domains. A graphical representation of this work is shown in Figure 1 and highlights how a variety of data can be modeled within PlasmoData.jl, which in turn provides access to diverse analysis tools and techniques.
@@
Figure 1: Overview of the PlasmoData.jl modeling framework. Diverse data objects can be modeled under the DataGraph abstraction; this provides access to a diverse analysis tools and techniques for analyzing the data
@@ 2 Graph Abstraction and Software Implementation
This section introduces the DataGraph modeling abstraction along with its implementation in PlasmoData.jl, available at https://github.com/zavalab/PlasmoData.jl. We introduce a mathematical formulation for the DataGraph and discuss how it can be applied to a variety of common data structures. We also provide brief code snippets showing how these representations are implemented in PlasmoData.jl. Versions of these code snippets are also available in the repository https://github.com/zavalab/JuliaBox/tree/master/PlasmoData_examples, along with the package and Julia versions used.
@@ 2.1 Mathematical Formulation
A DataGraph object is fully defined by its nodes, edges, adjacency matrix, ($A$ which encodes the graph connectivity), node data, edge data, graph data, node attributes, edge attributes, and graph attributes. Importantly, the adjacency matrix also encodes the type of edges in the graph (directed or undirected).
@@ 2.2 Software Implementation
Figure 2: Visualization of the DataGraph object. The PlasmoData.DataGraph (left) contains fundamental attributes (g, nodes, edges, node_map, edge_map, node_data, edge_data, graph_data) which correspond to attributes of the mathematical DataGraph object (right) containing nodes ($\mathcal{N}$), edges ($\mathcal{E}$), directionality ($A$), node data ($\bm{d}^{\mathcal{N}}$), edge data ($\bm{d}^{\mathcal{E}}$), graph data ($\bm{d}^{\mathcal{G}}$), and node, edge, and graph attributes ($\mathcal{A}_{\mathcal{N}}$, $\mathcal{A}_{\mathcal{E}}$, and $\mathcal{A}_{\mathcal{G}}$, respectively). Each node and edge can embed data of a variety of forms, such as scalar values, vectors, or other graphs
@@
PlasmoData.jl implements the DataGraph abstraction as shown above, and it provides a user-friendly interface and suite of functions; basic funcationality for constructing the graph is illustrated in Code Snippet 1. An empty PlasmoData.DataGraph is instantiated on Line 3. The data types within the {} define the types of data that can be defined in the node, edge, and graph data (see Figure 2). In this example, we use Any for the data type in the node, edge, and graph data which will be important when adding data to this graph. Alternatively, a user can call DataGraph() without the information in {}, and the default data types in the node, edge, and graph data will be Float64. Once the PlasmoData.DataGraph object is instantiated, nodes, edges, and data can be added to the graph. Nodes can be added to the PlasmoData.DataGraph by calling the add_node! function (Lines 6 - 10). PlasmoData.jl allows for any Julia object to be defined as a node. For example, as shown in the code snippet, the node names can be integers, strings, or symbols. This can be important, as will be shown later when representing matrices as PlasmoData.DataGraphs. Edges can be added to the PlasmoData.DataGraph by calling the add_edge! function (Lines 14 - 18). This function takes pairs of node names as arguments which represent the edge between those two nodes. These pairs of nodes can be passed as separate arguments (Lines 14 - 16) or as a tuple of two node names (Lines 17 - 18). This progressively builds the adjacency matrix of the graph object.
Code Snippet 1: Example showing the basic functionality for building a PlasmoData.DataGraph in PlasmoData.jl, including adding nodes, edges, and data and basic plotting functionality
1using PlasmoData, PlasmoDataPlots
@@
38PlasmoDataPlots.plot_graph(dg; xdim = 400, ydim = 400)
Figure 3: Visualization of the PlasmoData.DataGraph associated with Code Snippet 1
Data can also be embedded within the defined graph structure through API functions. Data is added to nodes through the add_node_data! function (Lines 21 - 25). This function takes the following arguments: i) the PlasmoData.DataGraph, ii) the node name, iii) the node data, and iv) the attribute name for the data (must be a string). The node data is restricted to be the type defined on the PlasmoData.DataGraph (in this case, the PlasmoData.DataGraph was instantiated with type Any, so any data can be passed). As seen in the code snippet, data have different formats, including floats, strings, vectors, matrices, or another PlasmoData.DataGraph. Data is added to edges through the add_edge_data! function (Lines 29 - 33). This function behaves similarly to the add_node_data! function, but in place of the node name argument, a pair of nodes (two separate arguments) or a tuple of two node names (one argument) is supplied. Data is added to the graph by the add_graph_data! function (Line 36). This function only takes the graph data and the graph attribute, as shown in the snippet. Additional node, edge, or graph data can be defined on the DataGraph by passing the data with a different attribute name. There is no limit to the number of attributes that can be added to the nodes, edges, or graph.
One of the benefits of graph representations is that they facilitate visualization of structure. For basic plotting capabilities, we have developed an accompanying package called PlasmoDataPlots.jl (available at https://github.com/zavalab/PlasmoDataPlots.jl) that interfaces to PlasmoData.jl and that enables visualizations of PlasmoData.DataGraph objects. This package provides basic functionalities for plotting both PlasmoData.DataGraphs and PlasmoData.DataDiGraphs primarily through the Julia packages Plots.jl and GraphMakie.jl. The visualization in Figure 3 is created by calling the PlasmoDataPlots.jl function plot_graph (Line 38).
@@ 2.3 Modeling Common Data Objects as DataGraphs
With the general mathematical formulation for the DataGraph defined, we next outline how specific data objects (matrices, 3D tensors, symmetric matrices) can be represented using the graph abstraction and how these can be implemented in PlasmoData.jl.Importantly, a benefit of representing matrices and tensors as graphs is that graphs are modeling objects that do not live in a Euclidean space. As such, all that matters from a graph perspective is connectivity (not location of nodes and edges in space). As such, a graph object is not affected by rotations (matrices are). This property is quite useful when processing images/video and also when processing data objects from molecular simulations (in which nodes move randomly in space) [34]. Thus the approaches in this section could be applied to a variety of problems in different fields.
@@ 2.3.1 Matrices as a Node-Weighted Graphs
We can represent a matrix $M\in\mathbb{R}^{p\times q}$ (where $p$ and $q$ are positive integers) as a node-weighted graph with a mesh structure [19, 35], and we can express the resulting graph using the DataGraph abstraction. We use the notation $m_{i,j}\in\mathbb{R}$ for the matrix entry at row $i$ and column $j$. For convenience, we define the integer set $\mathbb{N}_{k}:=\{1,2,...,k\}$. To represent the matrix as a graph, we define a node, $n_{*}$, for every matrix entry and place edges between all adjacent entries of the matrix. On each node, $m_{i,j}$ is stored as the node data. As there is a single weight in each node, we define the attribute set as $\mathcal{A}_{\mathcal{N}}=\{a\}$ (a singleton). Mathematically, this is represented with the DataGraph abstraction as:
@@
where $A$ is a symmetric matrix ($\mathcal{DG}$ is undirected) matching the connectivity defined by $\mathcal{E}$. An alternative structure can also be constructed by also adding edges between diagonal elements of the matrix [36]. In this case, the set of edges becomes =$\mathcal{E}=\{(n_{i,j},n_{i+1,j}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q}\}\cup\{(n_{i,j},n_{i,j+1}):i\in\mathbb{N}_{p},j\in\mathbb{N}_{q-1}\}\cup\{(n_{i,j},n_{i+1,j+1}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q-1}\}\cup\{(n_{i,j+1},n_{i+1,j}):i\in\mathbb{N}_{p-1},j\in\mathbb{N}_{q-1}\}$.
PlasmoData.jl facilitates the representation of matrices as graphs. Code Snippet 2 shows how a PlasmoData.DataGraph object can be constructed automatically by calling the matrix_to_graph function. Any matrix containing real numbers can be passed to this function. The boolean keyword argument diagonal can be passed to identify whether the diagonal edges should be included in the graph. The resulting graphs can also be visualized using PlasmoDataPlots.jl. In Line 11, nodes are given a fixed position (stored in the node data of the PlasmoData.DataGraph) for visualization. In Lines 14 - 21, the graph containing diagonals is plotted, where the weight data from the original matrix is used for coloring the nodes (see Line 19). Figure 4 visualizes the alternative representations. Because matrices themselves are a general abstraction for representing data, PlasmoData.DataGraph can be applied to many application. For instance, multivariate time series and grayscale images can be represented as matrices (and thus as graphs) [37, 38, 39].
Code Snippet 2: Example of representing a matrix as a PlasmoData.DataGraph
1using PlasmoData, Random, PlasmoDataPlots
@@
3Random.seed!(15)
4random_matrix = rand(12, 12)
5
6# Convert matrix to node-weighted graph
7matrix_graph_diags = matrix_to_graph(random_matrix; diagonal = true)
8matrix_graph_no_diags = matrix_to_graph(random_matrix; diagonal = false)
9
10# Fix the node positions
11set_matrix_node_positions!(matrix_graph_diags, random_matrix)
12
13# Plot the graph
14plot_graph(
@@
21)
Figure 4: Visualization of graph representations of a matrix. The top graph corresponds to the matrix_graph_no_diags PlasmoData.DataGraph object from Code Snippet 2, and the lower graph corresponds to the matrix_graph_diags PlasmoData.DataGraph object from Code Snippet 2 and is the output of the code snippet. Matrix entries are represented by nodes with the matrix entry values stored as data on each node
2.3.2 3D Tensors as Node-Weighted Graphs
Below we outline how a 3D tensor can be represented as a graph. While tensors are not restricted to be 3D, when we refer to tensors in this work, we will be referring to 3D tensors unless otherwise noted. Any tensor $T\in\mathbb{R}^{p\times q\times r}$ (where $p$, $q$, and $r$ are positive integers) can be represented as a node-weighted graph. There are at least a couple of methods for how to represent $T$. The first approach is to construct a node-weighted mesh following the description outlined in Section 2.3.1 based on the first couple of dimensions of $T$ (i.e., the mesh structure will be $p\times q$). The third dimension is then represented as a vector of weights attached to every element of the mesh (with $r=|\mathcal{A}_{\mathcal{N}}|$). This can be used, for instance, to represent color images (which have different color channels). In addition, we use the notation $t_{i,j,k}$ for the tensor entry, where $i,j,$ and $k$ are the entries in the first, second, and third dimensions, respectively. Under the DataGraph abstraction, this has the form:
@@
where $A$ is a symmetric matrix ($\mathcal{DG}$ is undirected) matching the connectivity defined by $\mathcal{E}$. While (7) is defined for 3D tensors, it could easily be extended to higher-order tensors.
PlasmoData.jl facilitates both of the representations discussed. In Code Snippet 3, the random tensor can be formed into the mesh structure discussed in Section 2.3.1 by calling the matrix_to_graph function (Line 7). This function recognizes that this is a 3D array and creates the mesh structure based on the first two dimensions. The third dimension are the weights on each node. The default name for these weights is the string "weight" with a number (e.g., "weight1", "weight2"), but a user can define their own names for the attributes in $\mathcal{A}_{\mathcal{N}}$. Alternatively, the user can call the function tensor_to_graph (Line 10) which performs the second method discussed above. These graphs can be visualized using PlasmoDataPlots.jl as before (Lines 13 - 20). In this case, the node locations are determined automatically within PlasmoDataPlots.jl through the NetworkLayout.jl package. A visualization of the methods in the code snippet is presented in Figure 5. Note that at this time, PlasmoData.jl only automates representing 3D tensors as graphs and not higher-order tensors.
Code Snippet 3: Example of representing a tensor as a PlasmoData.DataGraph using two different approaches
1using PlasmoData, Random, PlasmoDataPlots
@@
3Random.seed!(15)
4random_tensor = rand(4, 5, 6)
5
6# Convert the tensor to a node weighted graph with 4 x 5 nodes
7tensor_graph_2d = matrix_to_graph(random_tensor; diagonal = true)
8
9# Convert the tensor to a node weighted graph 4 x 5 x 6 nodes
10tensor_graph_3d = tensor_to_graph(random_tensor)
11
12# Plot the graph
13plot_graph(
@@
20)
Figure 5: Visualization of graph representations of a tensor. The upper graph corresponds to the tensor_graph_2d PlasmoData.DataGraph object from Code Snippet 3, and the lower graph corresponds to the tensor_graph_3d PlasmoData.DataGraph object from Code Snippet 3